Two boys enter running escalator at the ground floor in a shopping mall and they do some fun on it the first boy repeatedly follows P1 = one step up and Q1 equal to two step down where as the second boy repeatedly follows P2 equal to two Step Up and Q2 equal to one step down both of them move relative to escalator with speed Vr equal to 50 cm per second if the first boy takes t 1 equal to 250 seconds and second  boy takes t2 to equal to 50 seconds to reach the first floor how fast is the escalator running

Grade:12th pass

1 Answers

Arun
4 years ago
Let the speed of escalator be x.
speed*time=displacement
and displacement covered by both of them is from ground floor to first floor
therefore,
(50-x)250=(50+x)50.
12500-250x=2500+50x
-200x=-10000
x=(50cm/s)/2.
=25cm/s.